OpenStudy (anonymous):

I have $21.35 in change no pennies. If the # of quarters is 3 less than 4 times the # of nickles and the number of dimes is 14 more than the number of Nickles, find # each coin

OpenStudy (mertsj):

q=4n-3 d=n+14 5n+25(4n-3)+10(n+14)=2135 5n+100n-75+10n+140=2135 115n+65=2135 115n=2070 n=18 q=69 d=32